mathslover (mathslover):

Application of Derivatives... Problem Show that the function f(x) = 2x + sin x , x \(\in \) R is a strictly increasing function.

mathslover (mathslover):

@BSwan and @ganeshie8

mathslover (mathslover):

should I differentiate it?

OpenStudy (anonymous):

the definition :D

mathslover (mathslover):

f'(x) = 2 + cos x ( \(\forall x \in R\) )

OpenStudy (anonymous):

write it , and go through it

mathslover (mathslover):

Didn't get you? Did you mean to say to go through the definition?

OpenStudy (anonymous):

yep

mathslover (mathslover):

Well, this is what I think - -1 \(\le\) cos x \(\le\) 1 so, 1 \(\le\) 2 + \cos x \(\le\) 3 or 1 \(\le\) f'(x) \(\le\) 3

mathslover (mathslover):

And for the definition : A function f is said to be strictly increasing function in an interval R , iff \(x_1 < x_2\) \(\implies\) \(f(x_1) < f(x_2) \) \(\forall x_1 , x_2 \in R \)

OpenStudy (anonymous):

that tells nothing about strictly increasing , this tells u that the function is possitive and have upper/lowe bound

OpenStudy (anonymous):

dnt forget u have x

mathslover (mathslover):

And for the TEST of Monotonicity (for Strictly increasing functions : ) f(x) is said to be striclty increasing function in (a,b) iff f'(x) > 0 \(\forall\) x \(\in\) (a,b) .

OpenStudy (anonymous):

ic.. nw , u sued the differentiate

mathslover (mathslover):

Well, if I use the Test for Monotonicity , isn't it very clear that , as f'(x) > 0 (for all x belonging to R) ,so, the given function is strictly increasing... ?

OpenStudy (anonymous):

your correct !

mathslover (mathslover):

Oh well... Yeah! Thanks. Probably the easiest one I posted yet...!

OpenStudy (anonymous):

yep lol ur the master of Qn XD

mathslover (mathslover):

More to come... :)

OpenStudy (anonymous):

my Qn sadly no one answer them xD so feels jelouse

mathslover (mathslover):

I'm sorry for that.